OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x3 + 4 and g(x) = cuberoot(x-4)

OpenStudy (anonymous):

for example \(f(g(x))=(g(x))^3+4=(\sqrt[3]{x-4})^3+4=x-4+4=x \)

OpenStudy (anonymous):

u try to work on \(g(f(x))=x \)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw :)